OpenStudy (anonymous):

Simplify using basic identities cos of theta ^2 / sin of theta ^2 + csc θ sin θ

OpenStudy (anonymous):

22

OpenStudy (anonymous):

this ?\[\frac{ \cos^2(\theta) }{ \sin^2(\theta) } + \csc(\theta)\sin(\theta)\]

OpenStudy (anonymous):

\[= \cot^2(\theta) + \frac{ 1 }{ \sin(\theta) } * \sin(\theta) = \cot^2(\theta) + 1\] which i think is an identitity

OpenStudy (anonymous):

it is: \[\csc^2(\theta)\]